How do you solve: 6x+8/2?
Answer by checkley71(8403)

IF THIS EQUATION IS 6X+8/2=0 THEN 6X+4=0 OR 6X=-4 OR X=-4/6 OR X=-2/3
PROOF 6(-2/3)+8/2=0 OR -12/3+4=0 OR -4+4=0 OR 0=0